# Augmented doubly robust post-imputation inference for proteomic data

### v1

@@ 5 Peptide abundance in key proteins is associated with Alzheimer’s disease
Alz-heimer’s disease (AD) is a prominent neurodegenerative disorder among older adults. Numerous environmental and genetic factors are known to contribute to the disease, and related biological pathways have yet to be fully discovered. In this section, we apply the proposed method to identify important peptides associated with AD and related dementias. A bulk peptide-level dataset offers an opportunity to illustrate these methods in an important scientific setting (Merrihew et al., 2023). While samples in this dataset are annotated with a range of disease severity, we group them into two types; cases (samples with autosomal dominant/sporadic AD dementia) and controls (samples without dementia, with or without a high AD histopathologic burden). Two other covariates, the brain region, and PMI, are also used in our analysis.
@@
where we compute $p$-values associated with the type variable. BH procedure is then applied to convert them into q-values. The final discoveries are determined by applying a cutoff of 0.05 to them. Following this procedure, the Complete method selects 55, DR_W selects 50, DR_UW selects 58, and Plugin selects 79 peptides. The Complete, DR_W, and DR_UW methods have a similar number of discoveries. Only peptides with low missingness rates were recorded for these data, so it is reasonable that the number of discoveries for each method does not vary too much. Plugin selects more peptides than the other methods.
@@
MPLYGLHLWLPK sp—P03905— NU4M
Bhatia et al. (2022); Wesseling et al. (2017)
NLFTHLDDVSVLLQEIITEAR sp—O00408— PDE2A
Sheng et al. (2022); Shi et al. (2021); Delhaye et al. (2024)
TTHRPHPAASPSLK sp—Q01484— ANK2
Kumari et al. (2022); Higham et al. (2019)
MQNDTAENETTEKEEK sp—Q05682— CALD1
Pereira et al. (2017)
LGVSFLVLPK sp—Q16880— CGT
Tang et al. (2023); Moll et al. (2020); Ryckman et al. (2020)
NFEEFFNLHR sp—P20591— MX1
Prakash et al. (2024); Widjaya et al. (2023); Ma et al. (2012)
DVTHTC[+57]PSC[+57]K sp—Q9H305— CDIP1
Dileep et al. (2023); Inukai et al. (2021)
@@ 6 Discussion
In this paper, we present a statistical framework for analyzing proteomic data with missing values. Our proposed estimator $\hat{\bm{\beta}}_{UW}$ is established in a doubly robust framework and achieves reduced asymptotic variance leveraging correlations between different peptides. Through simulations, we demonstrate that the proposed estimator offers highly competitive statistical decisions in discovering signal peptides.
@@
In particular, we show that the proposed method possesses improved properties compared to other imputation-based methods, such as the Plugin method and the DR_W method. However, a final choice between imputation-based methods and the Complete method should depend on the quality of imputation achievable. Simulations and real data considered in Sections 4 and 5 provide a rich foundation for high-quality imputation with the VAE model. These data have Gaussian-like distributions, sufficient sample size, and robust correlation structure between peptides. If these conditions are not sufficiently satisfied, even the DR_UW method will not perform as well as the Complete method. For example, a simulation experiment reveals that if the outcome model in Section 4 produces a completely noisy imputation, then the doubly robust method will yield fewer discoveries compared to the complete method (Figure C7). However, even in such cases, the estimate $\hat{\bm{\beta}}_{UW}$ obtained by the DR method is similar to the estimate obtained by the Complete method (Figure C8). This follows because the consistency of $\hat{\bm{\beta}}_{UW}$ is still guaranteed by the doubly robust property, provided the estimated propensity score is consistent.
@@
The value of the proposed method depends on the quality of the imputation procedure. Although the approach is applicable regardless of the choice of imputation algorithm, we chose a refined VAE procedure that uses masking to robustly handle missing values as an integrated part of the procedure (Du et al., 2022). In the proteomic literature, one of the most commonly applied imputation procedures is a version of $k$-nearest neighbors (kNN). The standard kNN procedure in use involves imputing the missing values based on the mean of the $k$ closest peptides (Harris et al., 2023). Close peptides are used instead of close samples because it is very difficult to define close samples when there is excessive missingness. As a consequence, this kNN approach does not utilize low-dimensional covariates $\mathbf{W}_{i}$ in imputation and is potentially biased with fewer discoveries. An alternative approach is to adopt a two-step method: initially, we impute missing values based on the closest peptides and then, once the missing entries are filled, we impute the entire dataset based on the closest samples. When we apply this two-step approach to the AD dataset in Section 5, DR_W selects 65, DR_UW selects 51, and the Plugin method selects 121. Although we cannot reach an exact conclusion, the result of the VAE model in Section 5 is more aligned with theoretical expectations: DR_UW method provides more discoveries than the DR_W method.
@@
[Acknowledgments] We acknowledge Bernie Devlin for his help in contextualizing the findings in Section 5 within the scientific literature. The first author thanks Chan Park for assistance in identifying important reference papers.
@@
This project was funded by National Institute of Mental Health (NIMH) grant R01MH123184 and NSF DMS-2015492.
@@ B.3 Model training
Appendix C Supplementary figures/tables
Supplementary figures for Section 3.2
Figure C1: Simulation result of Model 1.
Figure C2: Simulation result of Model 2.
Figure C3: Simulation result of Model 4.
Supplementary figures and tables for Section 4
Figure C4: Histogram of correlation coefficient between peptides (left) and a 90% quantile absolute value of correlation coefficients computed for each peptide (right)
Figure C5: A result of realistic simulation with a single-cell dataset (Setting 1); FDR and TPR are summarized under different q-value cutoffs (0.01, 0.05 and 0.3)
Figure C6: A result of realistic simulation with a single-cell dataset (Setting 2); FDR and TPR are summarized under different q-value cutoffs (0.01, 0.05 and 0.3)
@@ Appendix C Supplementary figures/tables
Table C1: The proportions of peptides, whose corresponding proteins are included in protein lists corresponding to the peptides selected by the Complete method with a q-value cutoff of 0.05. Only the peptides additionally selected by each method compared to the Complete method (with a q-value cutoff of 0.01) are considered. A threshold 0.7 is applied to the observation rate of peptides.
Supplementary figures for Section 6
Figure C7: Volcano plot with a completely noisy imputation (a q-value cutoff=0.05)
Figure C8: Distribution of estimated coefficient $\hat{\beta}$ for the diameter variable under a completely noisy imputation

### v2

@@ 5 Peptide abundance in key proteins is associated with Alzheimer’s disease
Alzheimer’s disease (AD) is a prominent neurodegenerative disorder among older adults. Numerous environmental and genetic factors are known to contribute to the disease, and related biological pathways have yet to be fully discovered. In this section, we apply the proposed method to identify important peptides associated with AD and related dementias. A bulk peptide-level dataset offers an opportunity to illustrate these methods in an important scientific setting (Merrihew et al., 2023). While samples in this dataset are annotated with a range of disease severity, we group them into two types; cases (samples with autosomal dominant/sporadic AD dementia) and controls (samples without dementia, with or without a high AD histopathologic burden). Two other covariates, the brain region, and PMI, are also used in our analysis.
@@
where we compute p-values associated with the type variable. BH procedure is then applied to convert them into q-values. The final discoveries are determined by applying a cutoff of 0.05 to them. Following this procedure, the Complete method selects 55, DR_W selects 50, DR_UW selects 58, and Plugin selects 79 peptides. The Complete, DR_W, and DR_UW methods have a similar number of discoveries. Only peptides with a low missingness rate are shared for these data. Hence, there is limited opportunity for improvements over the complete method. While the Plugin method selects more peptides than the other methods, it is not surprising that the number of discoveries does not vary much across the well-calibrated methods.
@@
MPLYGLHLWLPK sp—P03905— NU4M Bhatia et al. (2022); Wesseling et al. (2017)
NLFTHLDDVSVLLQEIITEAR sp—O00408— PDE2A Sheng et al. (2022); Shi et al. (2021); Delhaye et al. (2024)
TTHRPHPAASPSLK sp—Q01484— ANK2 Kumari et al. (2022); Higham et al. (2019)
MQNDTAENETTEKEEK sp—Q05682— CALD1 Pereira et al. (2017)
LGVSFLVLPK sp—Q16880— CGT Tang et al. (2023); Moll, Shaw and Cooper-Knock (2020); Ryckman, Brockhausen and Walia (2020)
NFEEFFNLHR sp—P20591— MX1 Prakash et al. (2024); Widjaya et al. (2023); Ma et al. (2012)
DVTHTC[+57]PSC[+57]K sp—Q9H305— CDIP1 Dileep et al. (2023); Inukai et al. (2021)
@@ 6 Discussion
In this paper, we present a statistical framework for analyzing proteomic data with missing values. Our proposed estimator $\hat{\bm{\beta}}_{UW}$ is established in a doubly robust framework and achieves reduced asymptotic variance leveraging correlations between different peptides. Through simulations and real data analysis, we demonstrate that the proposed estimator offers highly competitive statistical decisions in discovering signal peptides.
The doubly robust method requires that the estimated conditional mean, $\hat{\nu}$, converges to ${\nu}$; however, it allows for this convergence to be slow—just not too slow. By contrast, under such circumstances, the natural competing imputation approach, the Plugin method, results in invalid inference (inflated FDR). At the performance extremes, there is nothing to gain: if $\hat{\nu}$ converges rapidly, then every imputation method works; and if $\hat{\nu}$ is inconsistent or converges very slowly, then none of the competing imputation methods are valid. This Goldilocks setting has two important implications: first, in all scenarios, the DR method never harms us, and there is a regime where it provides significant benefits; and second, we can roughly check the quality of the imputation through simulation results to determine the validity and efficiency of the DR approach.t‘
One might question why the doubly robust procedure offers significant advantages over simply replacing the missing data with the fitted mean value. Formally, let us consider two alternative estimators for the response $g(\cdot;\nu,\delta)$: plugin-missing, $\nu(\mathbf{W}_{i},\mathbf{U}_{i})+C_{i}(Y_{i}-\nu(\mathbf{W}_{i},\mathbf{U}_{i}))$, and doubly robust, $\nu(\mathbf{W}_{i},\mathbf{U}_{i})+\frac{C_{i}}{\delta(\mathbf{W}_{i})}(Y_{i}-\nu(\mathbf{W}_{i},\mathbf{U}_{i}))$. The plugin-missing estimator replaces $Y$ with our best estimate of $\nu$ only when the data are missing. For the doubly robust estimator, the first term also represents the best estimate of $\nu$, while the second term is designed to minimize the impact of the estimation error. Intuitively speaking, the $1/\delta$ term gives greater weight to the residuals in the portion of the data where we rarely observe $Y$. This adjustment has the effect of removing the first-order error in the estimation of $\nu$, as shown in Equation 3. This property was derived from the function $g$ satisfying $\mathbb{E}[\frac{\partial g(\cdot;\hat{\nu},\hat{\delta})}{\partial\hat{\nu}}]=\mathbb{E}[\frac{\partial g(\cdot;\hat{\nu},\hat{\delta})}{\partial\hat{\delta}}]=0$.
@@
In simulations, we show that the proposed DR method possesses improved properties compared to other imputation-based methods, such as the Plugin, Plugin-missing and DR_W methods. However, a final choice between imputation-based methods and the Complete method, which uses only observed values, should depend on the quality of imputation achievable. Simulations and real data considered in Sections 4 and 5 provide a rich foundation for high-quality imputation with the VAE model. These data have Gaussian-like distributions, sufficient sample size, and robust correlation structure between peptides. If these conditions are not sufficiently satisfied, even the DR_UW method will not perform as well as the Complete method. For example, a simulation experiment reveals that if the outcome model in Section 4 produces a completely noisy imputation, then the doubly robust method will yield fewer discoveries compared to the complete method (Figure C9). However, even in such cases, the estimate $\hat{\bm{\beta}}_{UW}$ obtained by the DR method is similar to the estimate obtained by the Complete method (Figure C10). This follows because the consistency of $\hat{\bm{\beta}}_{UW}$ is still guaranteed by the doubly robust property, provided the estimated propensity score is consistent.
@@
The value of the proposed method depends on the quality of the imputation procedure. Although the DR approach is applicable regardless of the choice of imputation algorithm, we chose a refined VAE procedure that uses masking to robustly handle missing values as an integrated part of the procedure (Du et al., 2022). In the proteomic literature, one of the most commonly applied imputation procedures is a version of $k$-nearest neighbors (kNN). The standard kNN procedure in use involves imputing the missing values based on the mean of the $k$ closest peptides (Harris et al., 2023). Close peptides are used instead of close samples because it is very difficult to define close samples when there is excessive missingness. As a consequence, this kNN approach does not utilize low-dimensional covariates $\mathbf{W}_{i}$ in imputation and is potentially biased with fewer discoveries. An alternative approach is to adopt a two-step method: initially, we impute missing values based on the closest peptides, and then, once the missing entries are filled, we impute the entire dataset based on the closest samples. When we apply this two-step approach to the AD dataset in Section 5, DR_W selects 65, DR_UW selects 51, and the Plugin method selects 121. Although we cannot reach an exact conclusion, the result of the VAE model in Section 5 is more aligned with theoretical expectations: the DR_UW method provides more discoveries than the DR_W method.
@@
[Acknowledgments] We thank the Editor and two anonymous reviewers for their helpful comments, Bernie Devlin for his assistance in contextualizing the findings in Section 5 within the scientific literature, and Chan Park for his help in identifying important reference papers. Haeun Moon is also affiliated with the School of Transdisciplinary Innovations (Integrative Data Science Program) and the Institute for Data Innovation in Science at Seoul National University.
@@
This project was funded by National Institute of Mental Health (NIMH) grant R01MH123184, NSF DMS-2015492 and DMS-2310764.
@@ B.3 Model training
Appendix C Supplementary figures and tables
Figure C1: Mean squared error of the three imputation methods in the setting of Model 3. The numbers are averaged over 200 repetitions (Section 3.2).
Figure C2: Simulation result of Model 1 (Section 3.2).
Figure C3: Simulation result of Model 2 (Section 3.2).
Figure C4: Simulation result of Model 3 under strong signals (Section 3.2).
Figure C5: Simulation result of Model 4 (Section 3.2).
Figure C6: Histogram of correlation coefficient between peptides (left) and a 90% quantile absolute value of correlation coefficients computed for each peptide (right) (Section 4)
Figure C7: A result of realistic simulation with a single-cell dataset (Setting 1); FDR and TPR are summarized under different q-value cutoffs (0.01, 0.05 and 0.3) (Section 4)
Figure C8: A result of realistic simulation with a single-cell dataset (Setting 2); FDR and TPR are summarized under different q-value cutoffs (0.01, 0.05 and 0.3) (Section 4)
@@ Appendix C Supplementary figures and tables
Table C1: The proportions of peptides, whose corresponding proteins are included in protein lists corresponding to the peptides selected by the Complete method with a q-value cutoff of 0.05. Only the peptides additionally selected by each method compared to the Complete method (with a q-value cutoff of 0.01) are considered. A threshold 0.7 is applied to the observation rate of peptides (Section 4).
Figure C9: Volcano plot with a completely noisy imputation (a q-value cutoff=0.05) (Section 6)
Figure C10: Distribution of estimated coefficient $\hat{\beta}$ for the diameter variable under a completely noisy imputation (Section 6)